# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

### v1

@@ Abstract
Genomic prediction of drug resistance in Mycobacterium tuberculosis remains challenging due to complex epistatic interactions and highly variable sequencing data quality. We present a novel Interpretable Variant-Aware Multi-Path Network (VAMP-Net) that addresses both challenges through complementary machine learning pathways. Path-1 employs a Set Attention Transformer processing permutation-invariant variant sets to capture epistatic interactions between genomic loci. Path-2 utilizes a 1D Convolutional Neural Network that analyzes Variant Call Format quality metrics to learn adaptive confidence scores. A fusion module combines both pathways for final resistance classification. We conduct comparative evaluations of unmasked versus padding-masked Set Attention Blocks, and demonstrate that our multi-path architecture achieves superior performance over baseline CNN and MLP models, with accuracy exceeding 95% and AUC of $\sim$97% for rifampicin (RIF) and rifabutin (RFB) resistance prediction. The framework provides dual-layer interpretability: Attention Weight Analysis reveals Epistatic networks, and Integrated Gradients (IG) was applied for critical resistance loci (notably rpoB), while gradient-based feature importance from the CNN pathway uncovers drug-specific dependencies on data quality metrics. This architecture advances clinical genomics by delivering state-of-the-art predictive performance alongside auditable interpretability at two distinct levels—genetic causality of mutation sets and technical confidence of sequencing evidence—establishing a new paradigm for robust, clinically-actionable resistance prediction.
@@ 1 Introduction
The rapid growth of high-throughput sequencing technologies has transformed genomics into a data-rich discipline, affording unprecedented opportunities to understand complex biological phenomena, such as genotype-phenotype associations and drug resistance mechanisms (sharma2022tuberculosis). Due to the contemporary requirement for personalized therapy within critical applications in Mycobacterium tuberculosis (MTB), drug resistance prediction from whole-genome sequencing (WGS) must be both accurate and interpretable. The prediction of drug resistance involves a number of significant modeling challenges. First, the genetic basis of drug resistance is determined by complex, high-order epistatic interactions among variants, meaning that the models need to learn set-based relationships rather than simple sequential dependencies. Second, raw WGS data from clinical samples are inherently of varying quality, with different coverages and allele frequencies, making it crucial to robustly distinguish the true pathogenic variants from low-confidence sequencing artifacts. In other words, the model should consider the reliability (confidence) of the input data in addition to the causality (biology) thereof. Overcoming the challenges of high-dimensional, variable-length genetic variant sequences and low-quality sequencing input is crucial for developing robust and biologically valid predictive systems (james2025whole; kim2022machine; cosgun2020exploring).
Recent advances in deep learning offer promising avenues for capturing the complex, non-local relationships inherent in genomic data (perelygin2025deep). Although CNNs have been widely applied to capture local patterns in genomic sequences, most often they require fixed-length representations and cannot model higher-order, long-range epistatic interactions among genetic variants that dictate complex phenomena such as drug resistance. This is partly overcome by Transformer-based architectures, which model non-linear, long-range dependencies using self-attention mechanisms with great success in sequence-based domains (wu2023transformer); however, applications to genomic variant sequences remain sparse. Moreover, existing approaches have fallen short, either treating the genome as a dense, image-like sequence (convolutional methods) or by being heavily reliant on expert-defined features, thereby limiting the discovery of novel epistatic relationships. Crucially, most current models treat critical VCF quality metrics as binary hard filters, lacking their treatment as continuous, learnable features that could modulate the model’s confidence in a variant’s biological contribution (eren2023improving).
In this work, we propose a novel Interpretable Multi-Path Fusion Architecture that strategically integrates a Set Attention Block (SAB) pathway with a dedicated, quality-aware 1D-CNN pathway to model drug resistance in MTB. We present a principled integration of a Set Transformer for learning permutation-invariant genetic causality and a 1D CNN for adaptive data quality assessment to overcome the limitations inherent in conventional sequential and feature-engineered models. The set of genomic variants (the reference-to-alternate change) is explicitly encoded with its precise position as a singular, biologically meaningful token (ChromPosi_Ref>AltChromPosi\_Ref>Alt) processed by the SAB pathway to capture complex, high-order epistatic variant-variant interactions under the principle of permutation invariance. The 1D-CNN pathway processes multi-channel representations of critical features in VCF format, (such as FRS and Genotype Confidence), enabling the model to learn an optimal weighting of data quality rather than relying on brittle, fixed thresholds for this task. A fusion module adaptively combines the outputs of these two pathways, leveraging both global, order-agnostic biological insights and local, quality-aware representations for superior predictive performance. This integrated approach extends the Set Transformer methods to genomic variant sequences while embedding a high degree of interpretability via the attention mechanism —a crucial requirement for clinical deep learning research.
Our work introduces the novel application of the Set Transformer architecture to analyze genetic variant sequences, culminating in a robust multi-path framework that significantly outperforms traditional MLP and CNN-only baselines by treating variants as an unordered set. We adapt the SAB in a two-fold manner: first, we establish a theoretically sound baseline by applying a purely permutation-invariant model, SAB without masking (lee2019set), and second, investigate how the same architecture behaves with the incorporation of padding masks while sustaining permutation invariance. Through extensive experiments and an in-depth interpretability analysis, we show how this architectural synergy enables the Set Transformer pathway to identify and map resistance loci accurately, while the parallel 1D-CNN provides a layer of confidence. Taken altogether, our contributions uniquely marry set-based attention modeling, multimodal fusion, and interpretability, bridging a critical gap between high-performance deep learning and actionable genomic modeling for precision medicine.
The rest of the paper is organized as follows. Section 2 summarizes the related work, including some deep learning frameworks and recent progress on deep genomic learning for drug resistance. Section 3 describes the dataset of genetic variants and VCF data used in our research. Section 4 describes the proposed method and strategy in detail, including the novel interpretable two-path transformer and 1D CNN framework. The experimental results and further discussion are presented in Section 5. Finally, Section 6 summarizes our contributions and perspectives on the possible future development of this research area.
@@ 2 Related Work
Genome-Wide Association Studies (GWAS) are powerful tools for identifying genetic variants, particularly single nucleotide polymorphisms (SNPs), linked to complex traits or diseases. However, traditional GWAS methods often face limitations in handling the complexity and scale of genomic data, especially in detecting non-linear interactions and epistatic relationships between SNPs. Recent advancements in machine learning (ML) and deep learning (DL) approaches have shown great promise in overcoming these limitations, offering improved accuracy and insights in post-GWAS analysis  (sun2021revisiting; sigala2023machine).
One of the primary advantages of ML over traditional statistical methods is its ability to identify complex interactions and hidden relationships between SNPs. However, traditional approaches such as linear regression (LR) are often inadequate in handling the high dimensionality and non-linearity of genomic data. Several works (an2020genome; maciukiewicz2018gwas; zhang2012identify) have used penalized regression techniques, like LASSO (L1) and Ridge regression (L2), to improve model interpretation and performance by shrinking coefficients, which helps in identifying significant SNPs with greater accuracy (sigala2023machine). But as mentioned previously and due to the Curse of Dimensionality problem, the computational complexity of applying these methods to genome-wide data was inevitable, consequently leading to the development of hybrid approaches that first pre-select SNPs using statistical methods before applying ML models. Random forests (RF) and gradient boosting machines (GBM) have been widely applied in genotype–phenotype association studies due to their ability to capture non‑linear relationships. For example,(arshadi2009predictive) showed that a GBM model was relatively robust to the inclusion of irrelevant SNPs, making it more tolerant to noise. Bayesian network methods have also been explored for modeling SNP–SNP and SNP–phenotype associations (sigala2023machine; kang2015inferring; ma2022sts). Support Vector Machines (SVMs) and Artificial Neural Networks (ANNs) are also used; while they show promise, benchmarking suggests they do not consistently outperform other methods, and their performance depends strongly on the size and quality of input data (nicholls2020reaching; sigala2023machine). For example, studies on GWAS‑case–control data have shown that encoding schemes (e.g., additive vs. homozygous/heterozygous) may affect classifier performance (mittag2015influence). Artificial Neural Networks have also been applied to SNP‑based classification, with some studies discussing how SNP encoding (e.g., homozygous/heterozygous representation) may affect performance (sigala2023machine). Deep learning models, such as the ExPecto framework (zhou2018deep) have demonstrated significant improvements in predicting the functional effects of SNPs across various tissues. ML methods are particularly effective in post-GWAS prioritization, where the goal is to narrow down the list of significant SNPs and genes identified during the GWAS stage. A variety of models, including logistic regression (LR), SVM, RF, GB, and ANN, have been applied to prioritize disease loci. (nicholls2020reaching) highlighted that random forests were top performers for diseases like amyotrophic lateral sclerosis and chronic kidney disease, emphasizing the importance of ensemble methods in improving prediction accuracy.
Other works adopt the approach of merging several techniques to enhance the performance of genotype–phenotype analysis and SNP prioritization. Combining machine learning models like random forests with traditional statistical methods has proven effective in improving prediction accuracy and scalability. (silva2022machine) demonstrated that random forest-based dimensionality reduction techniques can efficiently shrink the dataset size while preserving key genetic information. In their approach, random forest was used to rank SNPs according to their relevance, followed by clustering of the top-ranked SNPs for more focused analysis. Other studies combined RF (for SNP selection) and used KNN and SVM algorithms to classify the identified SNPs for their association with asthma (gaudillo2019machine). Studies in (boutorh2016complex; boutorh2015classication; boutorh2014grammatical) focused on dimensionality reduction in the genomic context, where their hybrid framework based on association rule mining (ARM), grammatical evolution, and neural networks addressed the “large p, small n” challenge (i.e., large features, relatively few samples) by selecting informative SNP‐sets and reducing feature space prior to classification.
The application of deep learning (DL) models presents significant advancements, particularly in overcoming the limitations associated with ML methods and traditional statistical methods, such as linear models, that struggle with high-dimensional genomic data. One of the primary advantages of DL models (especially CNNs) is their capacity to model non-linear interactions between SNPs, as well as their ability to incorporate epistatic relationships, which can enhance the accuracy of disease association predictions. For instance, (liu2018chromatin) introduced DEOPEN, an artificial neural network that combines a deep convolutional network with a three-layer feed-forward neural network to predict chromatin accessibility and evaluate genetic variants in breast cancer. Authors in (liu2019phenotype) further developed an independent deep CNN model for predicting phenotypic traits from SNP data, employing a saliency map visualization method to identify significant biomarkers and demonstrating improved accuracy over traditional statistical methods. Similarly, (romagnoni2019comparative) utilized dense neural networks with various architectures, achieving competitive mean AUC scores around 0.80 in classification tasks. (wang2019deep) introduced a deep mixed model that integrates CNN and Long Short-Term Memory (LSTM) architectures. Another key example of this is the ExPecto framework developed by (zhou2018deep), which utilizes deep convolutional neural networks, L2-regularized linear models, and spatial feature transformation techniques for in silico prediction of gene expression and disease risk. Similarly, DeepSEA (zhou2015predicting) offers a deep learning framework that uses convolutional layers to predict the effects of noncoding variants by learning regulatory sequences from chromatin-profiling data. It integrates various chromatin information to capture complex interactions, enabling precise single-nucleotide predictions and aiding in prioritizing functional variants like eQTLs and disease-associated variants. DeepWAS (arloth2020deepwas) integrates deep learning predictions of regulatory effects into a multivariate GWAS framework, addressing the challenge of linking genetic variants to their functional roles. By annotating SNPs for regulatory potential and using regularized regression models, DeepWAS reduces the multiple testing burden and incorporates current regulatory predictions. This method has been successfully applied to GWAS data for conditions like multiple sclerosis and major depressive disorder, effectively identifying functionally relevant SNPs. Another model is DeepCombi (mieth2021deepcombi) which addresses the issue of missing heritability in traditional GWAS methods by combining convolutional neural networks (CNNs) with GWAS analyses to capture complex genetic interactions and multi-locus effects. Others like (kwon2022genome) also utilized CNNs to develop CNN-GWAS that focus on enhancing the early diagnosis of atrial fibrillation (AF). Another example is the GenNet framework (van2021gennet) designed to tackle the complexities of interpreting genetic data in GWAS while addressing computational challenges. It managed to outperform traditional methods, achieving significant accuracy and identifying both well-replicated and novel genes linked to traits such as hair color and schizophrenia.
Transformer-based architectures are also widely used with genomic data, particularly for their ability to handle large sequence data, which is often the case with long DNA sequences. One of the leading models is DNABERT, introduced in (ji2021dnabert), a pre-trained Bidirectional Encoder consisting of 12 transformer blocks, designed to capture k-mer representations of DNA for downstream tasks such as promoter and enhancer recognition. Another work introduced LOGO (yang2022integrating), a lightweight language model that combines self-attention and convolutional layers to efficiently interpret non-coding regions of the human genome at high resolution, demonstrating improved performance with fewer parameters. More recent developments have extended these approaches by integrating hybrid convolution–transformer designs, hierarchical attention mechanisms, and domain-specific pretraining strategies to better capture long-range dependencies while maintaining computational efficiency.
Research in (green2022convolutional) trained convolutional neural networks on variants from 18 genomic loci across $10\,\text{k}$ M. tuberculosis isolates to predict resistance to 13 drugs; their multi-drug CNN (MD-CNN) and single-drug CNNs achieved high AUCs and, using saliency mapping, recovered known resistance sites while nominating novel candidate positions for further validation. (kuang2022accurate) compared traditional ML classifiers and CNNs on whole-genome sequence–derived features to build a fast, practical pipeline for calling TB drug resistance; they showed that carefully tuned classical models can match deep models on many tasks and presented workflows optimized for clinical turnaround. GWAS-derived SNP sets and deep learning (CNN) were used in (kwon2022genome) to predict phenotype (atrial fibrillation) from genotype, demonstrating how GWAS hits can be repurposed as structured inputs for predictive models and highlighting pitfalls and opportunities of transferring association signals into individual-level predictors. A recent research work (saliba2025enhanced) introduced group-association modeling to jointly model correlated resistance phenotypes across pathogens, leveraging shared patterns to boost sensitivity for multi-drug resistance detection while keeping model complexity low. (deelder2019machine) present flexible machine-learning frameworks that take large numbers of WGS-derived predictors to classify resistant vs susceptible M. tuberculosis isolates; their results demonstrate robust predictive performance across drugs and emphasize scalable feature handling and interpretability for clinical usage . (groschel2021gentb) deliver GenTB, a user-facing ML pipeline and web tool that integrates WGS inputs with trained models to predict TB resistance profiles; they benchmark the system against other tools, prioritize usability for labs, and provide transparent outputs to aid interpretation in clinical workflows . TB-DROP (wang2024tb), an end-to-end deep learning model that consumes whole-genome mutation matrices to predict drug resistance and provide a lightweight GUI; the work focuses on clinical applicability with fast per-sample runtime and validate performance on diverse cohorts. An MDPI Diagnostics study systematically evaluates multiple ML strategies on mutation-based inputs for resistance prediction, comparing feature engineering pipelines and classifiers, and discussing limits in generalization that underscore the need for diverse training cohorts and robust external validation (paredes2025predicting).
In summary, recent advancements in genomic data analysis reflect a clear evolution from traditional GWAS and classical machine learning methods toward more sophisticated hybrid and deep learning architectures. These modern approaches increasingly integrate statistical modeling, feature selection, and end-to-end representation learning to capture the complex, non-linear relationships inherent in genomic data. Despite these advances, several key challenges remain unresolved—particularly those concerning model interpretability, cross-population generalizability, and the scalability of learning algorithms to high-dimensional, heterogeneous variant data. Addressing these challenges requires the development of architectures capable of unifying biological interpretability with data-driven representation power. The following sections describe the dataset and our proposed interpretable multi-path deep learning framework, designed specifically to meet these needs.
3 Data Description

### v2

# VAMP-Net: An Interpretable Multi-Path Network of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

@@ Abstract
Genomic prediction of drug resistance in Mycobacterium tuberculosis is often hindered by complex epistatic interactions and variable sequencing quality. We present the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a novel architecture addressing these challenges through a dual-pathway approach. Path-1 utilizes a Set Attention Transformer to model permutation-invariant variant sets and capture epistatic dependencies, while Path-2 employs a 1D-CNN to analyze VCF quality metrics for adaptive confidence scoring.Evaluated on four critical anti-TB drugs, VAMP-Net significantly outperforms baseline CNN and MLP models, achieving accuracies >> 95% and AUCs $\sim$0.97 for Rifampicin and Rifabutin. Feature attribution analysis via Integrated Gradients successfully recovered canonical targets (rpoB, embB, katG) and discovered high-impact novel loci. Functional enrichment confirmed these novel variants constitute non-random metabolic modules (p=0.00239) centered on cell-wall remodeling.Furthermore, systematic ablation of the Quality-Aware pathway demonstrates that the model performs a learned "integrated audit," prioritizing the Fraction of Supporting Reads and relative confidence over raw depth to mitigate technical noise. This dual-layer interpretability, bridging genomic pathogenicity with technical reliability, establishes a new paradigm for robust, auditable, and clinically actionable resistance prediction, positioning VAMP-Net as an important tool for both diagnostic classification and mechanistic discovery in clinical genomics.
@@ 1 Introduction
The rapid growth of high-throughput sequencing has transformed genomics into a data-rich discipline, offering unprecedented opportunities to decode complex biological phenomena like drug resistance in Mycobacterium tuberculosis (MTB) [39]. For clinical applications, drug resistance prediction from whole-genome sequencing (WGS) must be both accurate and interpretable. However, modeling MTB resistance presents two primary challenges: (1) the genetic basis is governed by high-order epistatic interactions, requiring models that learn set-based relationships rather than simple sequential dependencies; and (2) clinical WGS data varies significantly in quality, making it crucial to distinguish true pathogenic variants from sequencing artifacts. Effectively balancing biological causality with data reliability is essential for developing robust predictive systems [20, 23, 11].
Recent deep learning advances, particularly Convolutional Neural Networks (CNNs) and Transformers, offer promising avenues for capturing non-local relationships in genomic data [32, 47]. While CNNs excel at local pattern recognition, they often require fixed-length inputs and struggle with the long-range epistatic interactions inherent in drug resistance. Conversely, while Transformers can model these dependencies, their application to unordered sets of genomic variants remains sparse. Furthermore, existing models typically treat critical variant calling (VCF) quality metrics as binary hard filters, discarding potentially valuable continuous information that could modulate the model’s confidence in a variant’s biological contribution [13].
Despite recent advancements, existing models frequently treat genomic variants as rigid sequences and rely on static, "hard" VCF filters. This approach is limited to account for the fundamentally permutation-invariant nature of genomic variant sets and the continuous spectrum of sequencing quality metrics. Current methods either ignore technical noise entirely or rely on brittle, expert-defined thresholds that cannot capture the nuanced relationship between data confidence and predictive certainty.
To address these gaps, we propose VAMP-Net, an interpretable multi-path fusion architecture designed to bridge the gap between genomic deep learning and clinical microbiology. Our key contributions are summarized as follows:
  • Expert-Informed Multi-Path Fusion: We introduce a dual-stream architecture that decouples biological signals from technical noise. By integrating a Set Attention Block (SAB) pathway with a quality-aware 1D-CNN, VAMP-Net mirrors the "clinical audit" workflow of expert microbiologists, allowing technical confidence to dynamically modulate biological predictions through a Gated Adaptive Fusion mechanism.
  • Permutation-Invariant Modeling via SAB: Unlike sequence-based Transformers, our model utilizes the Set Attention Block (SAB) to capture high-order epistatic interactions. This approach enforces a permutation-invariant inductive bias, ensuring the model respects the biological topology of the isolate as an unordered set of variants rather than an artificial sequence.
  • Clinically-Grounded Interpretability: We provide a validation framework that maps Path-1 attention weights directly to established resistance hotspots (e.g., katG, rpoB). By demonstrating that high-attention regions align with verified microbiology, we confirm that VAMP-Net’s decision-making is driven by biological causality rather than spurious statistical artifacts.
The rest of the paper is organized as follows. Section 2 summarizes related work. Section 3 details the proposed Multi-Path Network. Section 4 presents the experimental results, interpretability analysis, and discussion. Section 5 summarizes our contributions and future perspectives.
@@ 2 Related Work
The evolution of drug resistance prediction in M. tuberculosis (MTB) has progressed from simple association studies to complex deep learning frameworks. This section categorizes prior efforts and identifies the specific gaps addressed by the proposed Multi-Path architecture, as summarized in the comparative overview in Table 1.
2.1 Statistical and Classical Machine Learning
Genome-Wide Association Studies (GWAS) remain the foundation for identifying single nucleotide polymorphisms (SNPs) linked to drug resistance [39, 44]. Traditional methods, however, struggle with high-dimensional data and non-linear epistatic interactions [41]. To overcome these, penalized regression techniques such as LASSO and Ridge have been employed to identify significant loci [1, 28]. Ensemble methods, including Random Forests (RF) and Gradient Boosting Machines (GBM), have demonstrated robustness to genomic noise [29, 27]. Hybrid approaches have further enhanced performance by using RF for dimensionality reduction followed by clustering or Support Vector Machines (SVM) for classification [42, 14]. Prior studies in [5, 4, 3] focused on dimensionality reduction through hybrid frameworks based on association rule mining (ARM) and ANN to address the “large $p$, small nn” challenge. More recently, [12] and [38] introduced flexible machine-learning and group-association frameworks to jointly model correlated resistance phenotypes, while systematic evaluations of ML strategies on mutation-based inputs have highlighted the utility of standardized feature sets [31]. Despite their utility, these models often rely on expert-defined feature engineering and treat sequencing quality metrics as binary filters rather than informative features.
2.2 Deep Learning and Monolithic Architectures
Deep learning (DL) has significantly advanced post-GWAS prioritization by capturing non-linear interactions through architectures like DeepSEA and ExPecto [53, 52]. Convolutional Neural Networks (CNNs) have become a standard for predicting phenotypic traits from SNP data due to their ability to model local patterns [26, 35]. Specifically for MTB, models like MD-CNN and CNN-GWAS have successfully integrated CNNs with genomic analysis to address missing heritability [16, 24]. Other frameworks have utilized monolithic deep structures, such as DeepAMR [51], which employed denoising autoencoders to learn latent representations of co-occurrent resistance, and MD-WDNN [7], which pioneered a wide-and-deep approach to share information across multiple drugs. Furthermore, GenTB [17] and TB-DROP [46] provided fast, clinically-applicable resistance calls using multi-task and MLP-based representations. While successful, these models often rely on fixed-length mutation vectors that struggle with the “long-tail” of rare variants and show limitations to distinguish between biological signals and technical sequencing noise.
2.3 Transformer-based and Graph Modeling
The emergence of Transformer architectures, such as DNABERT [22], has enabled the processing of large DNA sequences through self-attention mechanisms. Lightweight models like LOGO combine self-attention and convolutional layers to interpret non-coding regions with high resolution [49]. Modern sequence models like LLMTB [45] leverage BERT-based tokens to capture complex genomic patterns. This shift toward attention-driven models is reflected across clinical diagnostics, from cardiovascular disease detection [30] to oncology-driven feature extraction [40]. Additionally, graph-based approaches have been introduced to model relational dependencies; for example, HGAT [50] employs heterogeneous graph attention networks to model the topological relationships between isolates and mutations, while [6] utilized graph representation learning for drug repurposing.
2.4 The VAMP-Net Position: Expert-Informed Decoupling
Despite the sophistication of current state-of-the-art (SOTA) models, a significant gap remains: existing architectures typically treat genomic data as a uniform black box, limited to isolate biological causality from technical artifacts. As illustrated in Table 1, VAMP-Net addresses this critical limitation through an expert-informed multi-path architecture. Unlike the sequence-dependent logic of LLMTB or the fixed-graph constraints of HGAT, VAMP-Net utilizes a permutation-invariant Set Attention Transformer (Path-1) to capture non-linear epistatic interactions as unordered biological sets. Crucially, it introduces a dedicated Technical Auditor (Path-2) using a 1D-CNN to process quality-aware metrics, a feature absent in the monolithic designs of DeepAMR or DeepPTB. By integrating these streams via an Adaptive Fusion mechanism, VAMP-Net mirrors the clinical audit-trail logic used by microbiologists, providing a level of diagnostic reliability and auditable interpretability that current models struggle to achieve.
Table 1: Comparison of architectural properties and training datasets across prior genomic modeling approaches versus VAMP-Net.
Method Model Type Variant Repr. Quality Handling Interpretability Dataset Used
WHO/GWAS Statistical Categorical SNPs Hard Thresholds High (Manual) CRyPTIC ($N\approx 10k$)
GenNet CNN Hierarchical Image Binary Filter Feature Map Public Repos ($N\approx 20k$)
DeepAMR Autoencoder Binary Matrix None Latent Space Global Cohort ($N\approx 16k$)
MD-WDNN Wide & Deep Fixed-Length Vector Static Input Weights PATRIC/ReSeqTB ($N\approx 3.6k$)
HGAT GNN Heterogeneous Graph Implicit Edge Importance Kaggle/AMK ($N\approx 4.4k$)
DNABERT Transformer k-mer Sequence N/A Self-Attention Ref. Genomes
LLMTB Transformer BERT Tokens Sequence-only Attention Maps CRyPTIC ($N\approx 12k$)
TB-DROP MLP/CNN Mutation Matrix Fixed Filters Low Curated Repos ($N\approx 4k$)
VAMP-Net Multi-Path Unordered Set Adaptive Fusion Gated Logic CRyPTIC ($N\approx 10k$)
3 Methodology
Our approach introduces a Variant-Aware Multi-Path Network (VAMP-Net) that synergistically couples a Set Transformer architecture with a quality-aware 1D-Convolutional Neural Network for genetic variant analysis. This network is designed as a two-branch fusion model presented in Figure 1.
3.1 VAMP-Net Overview:
A).Path-1: The first branch, a Set Transformer, is specifically engineered to process the genetic variant sequence as an unordered set of tokens. This branch employs the Set Attention Block (SAB) [25] to capture the complex non-sequential relationships between the variants, reflecting the inherent permutation-invariant nature of a set.
B). Path-2: The second branch of our framework is a quality-aware 1D-CNN. This path focuses on local sequential features. For each variant, it takes as input a corresponding quality score sequence. The 1D-CNN is efficient in capturing the pattern in these local quality signals, thus offering the model critical information on the confidence and reliability of a variant call.
Figure 1: Overview of the Variant-Aware Multi-Path Network (VAMP-Net) Architecture. The framework employs a dual-pathway design to disentangle biological signal from technical noise. Path 1 (Biological Topology) utilizes a Set Attention Transformer (SAB) to process genomic variant sequences as unordered sets; the multi-head attention mechanism captures high-order epistatic interactions between mutations while maintaining permutation invariance. Simultaneously, Path 2 (Technical Confidence) employs a Quality-Aware 1D-CNN to process multi-channel VCF features, extracting local patterns of sequencing reliability. The outputs of both paths are integrated at a Late Fusion Node via an adaptive gating mechanism, allowing technical confidence to modulate biological findings. The fused representation is finally processed by a Classification Head consisting of fully connected layers and dropout regularization to produce the final drug resistance prediction.
C). Fusion Module: Finally, the outputs of the Set Transformer and the 1D-CNN branches are fused through a late-fusion mechanism, as pesented in the Figure 1. The representations from both branches are concatenated and passed to a final classification layer. This fusion allows the model to make predictions based on a rich, multi-faceted understanding of the data: the global, set-based relationships between variants from the SAB, and the local, sequential quality signals from the 1D-CNN. This multi-scale learning (local + global) provides both robustness (resistance to noisy data) and biological interpretability (attention highlighting key variants).
To formulate the problem addressed, a genomic sample is represented as a set of $N$ genetic variants: $V=\{v_{1},v_{2},...,v_{N}\}$ where each variant $v_{i}$ is described by $(\textbf{g}_{i},\textbf{f}_{i})$, with $\textbf{g}_{i}$ representing a sequence of genetic alteration and $\textbf{f}_{i}\in\mathbb{R}^{d}$ denoting the associated feature vector extracted from Variant Call Format (VCF) files. The prediction task is binary classification: $y=f_{\theta}(V)$, where $y\in\{0,1\}$ indicates resistance or susceptibility to a given drug, and $f_{\theta}$ is a parameterized deep learning model. The VAMP-Net approach is presented in Algorithm 1.
Algorithm 1 VAMP-Net Multi-Path Inference and Adaptive Fusion, performed on an NVIDIA H100 GPU.
1:Set of Genomic Variants $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, VCF Quality Metrics $\mathbf{Q}\in\mathbb{R}^{n\times d}$
2:Resistance Probability $\hat{y}\in[0,1]$
3:
4:// Path 1: Biological Signal Extraction
[…]
7:
8:// Path 2: Technical Context Assessment
[…]
10:
11:// Adaptive Fusion and Gating
[…]
13:if Mechanism is Suppression then
14:  $\mathbf{Z}_{fused}\leftarrow\mathbf{Z}_{bio}\odot\mathbf{g}$ $\triangleright$ Down-weight low-confidence signals
15:else if Mechanism is Amplification then
16:  $\mathbf{Z}_{fused}\leftarrow\mathbf{Z}_{bio}+(\mathbf{Z}_{bio}\odot\mathbf{g})$ $\triangleright$ Enhance high-confidence signals
17:end if
18:
19:// Final Prediction
[…]
22:return $\hat{y}$
3.2 Expert-Informed Design: The Rationale for Functional Decoupling
The choice of these two specific architectures is a mathematical and biological necessity based on the different "shapes" of the genomic data. Genomic data in a clinical setting is heterogeneous. It contains two entirely different types of information that should not be processed the same way.
1. Why Permutation-Invariant Set Attention? (The Biology):
  • The Nature of Mutations: Genomic variants in an isolate function as a biological set rather than a linear sequence. Whether a mutation in katG appears before or after a mutation in rpoB in the data file is irrelevant to the bacteria’s resistance.
  • Inductive Bias: Standard Transformers or RNNs assume a sequence (like a sentence). By using Set Attention (SAB) and removing positional encodings, we enforce a "set" inductive bias. This tells the model: "Focus on which mutations are present together (epistasis), not the order they were recorded."
  • Capturing Epistasis: Attention mechanisms are mathematically superior at capturing non-linear interactions between distant variants (e.g., how a compensatory mutation in ahpC interacts with a primary mutation in katG).
2. Why 1D-CNN? (The Technical Auditor:
  • The Nature of Sequencing Quality: VCF features (Depth, Fraction of Supporting Reads, Confidence scores) are local, quantitative signals.
  • Pattern Extraction: 1D-CNNs are the gold standard for extracting features from local neighborhoods of multi-channel data. They are designed to detect "shapes" in the noise, such as a specific drop in read depth across a variant, that signify a low-quality call.
  • Feature Differentiation: If we fed these technical metrics into the Transformer along with the variants, the high-magnitude "quality" numbers might overwhelm the sparse "biological" signals. The 1D-CNN acts as a pre-processor that compresses raw technical noise into a single "Confidence Vector."
3. Biomimetic Rationale: The Clinical Audit Workflow:
Beyond mathematical necessity, the dual-pathway architecture of VAMP-Net is explicitly designed to mirror the diagnostic reasoning of expert clinical microbiologists. In a manual diagnostic setting, an expert does not evaluate a mutation in isolation; instead, they engage in a "multi-path" heuristic:
  • Biological Assessment (Path-1): The expert identifies the presence of specific genetic markers (e.g., katG or rpoB mutations) and evaluates their known association with drug resistance.
  • Technical Auditing (Path-2): Simultaneously, the expert scrutinizes the raw sequencing run’s quality, verifying that the read depth and genotype confidence are sufficient to support a clinical call.
By formalizing these distinct tasks into specialized architectural branches, VAMP-Net avoids the "black-box" pitfalls of monolithic models. This functional decoupling ensures that the final prediction is not merely a statistical correlation, but a clinical decision where the biological signal is rigorously audited by technical reliability, a process we term Expert-Informed Fusion.
3.3 Dataset for MTB Drug Resistance:
The dataset used in this study comprises genetic variant sequences of Mycobacterium tuberculosis (MTB). Each sample is represented by a set of all identified genetic variants, resulting in sequences of variable length.
3.3.1 Data Description